# A Survey on Human-Centered Evaluation of Explainable AI Methods in Clinical Decision Support Systems

### v1

@@ Abstract
Explainable AI (XAI) has become a crucial component of Clinical Decision Support Systems (CDSS) to enhance transparency, trust, and clinical adoption. However, while many XAI methods have been proposed, their effectiveness in real-world medical settings remains underexplored. This paper provides a survey of human-centered evaluations of Explainable AI methods in Clinical Decision Support Systems. By categorizing existing works based on XAI methodologies, evaluation frameworks, and clinical adoption challenges, we offer a structured understanding of the landscape. Our findings reveal key challenges in the integration of XAI into healthcare workflows and propose a structured framework to align the evaluation methods of XAI with the clinical needs of stakeholders.
@@ 1 Introduction
Artificial Intelligence (AI) is increasingly being integrated into Clinical Decision Support Systems (CDSS) to improve timely diagnosis, optimize treatment plans, and reduce healthcare costs Sendak et al. (2020). Estimates suggest that AI adoption in healthcare could reduce U.S. healthcare expenditures by 5-10%, translating to \$200–\$360 billion in annual savings while maintaining quality and access Sahni et al. (2023). The rapid growth of AI-driven CDSS has also attracted significant investment, with venture capital funding in AI in health exceeding \$11 billion in 2024 Goldsack and Overgaard (2024). These trends indicate that the role of AI in medical decision making will continue to expand.
However, a key challenge accompanying this expansion is the necessity for trust and transparency in AI-driven medical decisions. Despite advancements in deep learning and other black-box AI models, clinicians often hesitate to adopt AI-generated recommendations due to their lack of interpretability Hou et al. (2024). Trust in CDSS is critical because errors in automated decision-making can have life-threatening consequences Zhang and Zhang (2023). In response, Explainable AI (XAI) has emerged as a solution to improve model transparency, inderectly allowing medical professionals to better understand why an AI system makes certain predictions Loh et al. (2022). Additionally, regulatory frameworks, such as the General Data Protection Regulation (GDPR), emphasize the need for explainability in automated decision-making, thereby necessitating interpretable AI solutions in healthcare.
@@
To address this gap, this paper presents a systematic survey of human-centered evaluations of XAI methods in CDSS, focusing on the intersection of explainability, usability, and clinical decision-making. Specifically, we: (1) categorize existing works based on XAI methodologies, evaluation frameworks, and clinical adoption challenges, offering a structured synthesis of the field; (2) identify key gaps in current evaluation practices from clinicians perceptions of XAI-based CDSS, revealing that while XAI is generally well-received, significant challenges remain—such as high cognitive load, misalignment with clinical knowledge, and lack of early involvement with stakeholders; (3) introduce a conceptual framework to bridge the socio-technical gap in AI-driven CDSS by aligning XAI evaluation methods with the needs of diverse stakeholders (e.g., clinicians, hospital administrators, AI developers); (4) propose a medical field-agnostic, stakeholder-centric approach for CDSS development, advocating that CDSS should be designed as human-augmentation tools rather than replacements.
To theoretically ground this misalignment between human expectations and technological capabilities, we leverage Ackerman’s socio-technical gap theory, which highlights the challenges arising from such discrepancies Ackerman (2000). This analysis informs our proposed framework, emphasizing stakeholder-driven evaluations to ensure CDSS solutions are clinically relevant, interpretable, and effectively integrated into healthcare workflows. By aligning technical advancements in XAI with rigorous human-centered evaluations, we aim to steer research toward developing trustworthy, deployable CDSS solutions.
@@ 2.1 Clinical Decision Support Systems (CDSS)
CDSS have long been integral to medical decision-making, assisting clinicians in enhancing patient outcomes. Initially emerging in the late 1950s with rule-based systems, CDSS evolved significantly with advancements in AI. Modern CDSS now leverage sophisticated machine learning algorithms to process vast datasets efficiently. Concurrently, the rise of XAI has enabled the integration of interpretable solutions, essential for fostering trust and promoting the adoption of CDSS in clinical practice Ghassemi et al. (2021).
Operationally, CDSS utilize patient data to generate specific treatment recommendations Musen et al. (2021). The back-end employs machine learning algorithms trained for specific tasks, processing data from Electronic Health Records (EHRs). EHRs include both structured data (e.g., tabular data) and/or unstructured data (e.g., text, images), representing comprehensive digital versions of patients’ medical histories, including diagnoses, treatments, and test results. System performance is evaluated using machine learning metrics pertinent to the specific task. Synchronously, the front-end runs a user interface (UI) connected to the back-end that clinical practitioners are supposed to interact with. Such UI displays patient data, alerts, recommendations (sometimes with explanations), and decision-support reports in an accessible manner to facilitate clinical decision-making. To this end, human-centered evaluations are conducted to assess the overall effectiveness and efficiency of the system in the specific task. However, researchers have identified that the absence of such evaluations is a primary factor in the lack of adoption of AI-based CDSS solutions  Musen et al. (2021).
In summary, the integration of XAI together with human-centered evaluation assures that the system is usable within clinical workflows and more transparently meets the needs of stakeholders to foster acceptance of CDSS.
2.2 Explainability and XAI Methods: a Taxonomy
From a psychological standpoint, humans seek explanations to predict future events, as explanations facilitate generalization Vasil and Lombrozo (2022). Unlike descriptions, explanations provide understanding by identifying “difference-makers” in causal relationships. In medicine, where transparency and accountability are crucial, explanations may offer insights into the rationale behind AI-based CDSS recommendations, enabling practitioners to align the model’s reasoning with their medical expertise.
In AI, explainability and interpretability improve people’s understanding of such models, driving from suboptimal to “better” outcomes Boyd-Graber et al. (2022). More in general, the main principle of XAI is to make the decision-making process of AI systems transparent and understandable to humans. Many taxonomies have been proposed to categorize XAI methodologies (e.g., see Speith (2022)). In this paper, we adopt the conceptual approach based on the following macro criteria (for a more comprehensive review, please refer to Molnar (2022)):
@@ 2.2.4 Output: Numerical or Visual
Visual output methods are particularly suitable for deep learning models operating with unstructured data such as images. Visualization techniques, such as saliency maps, class activation maps (CAM), and Grad-CAM, enable the interpretation of feature importance in CNNs, highlighting regions of interest in input images. Parallely, this also extends to token importance visualisation for language models. Beyond deep learning, structured data can also benefit from visual representations, including partial dependence plots, feature importance bar charts, and decision boundary plots. However, we treat those methodologies as numerical. Methods: all methodologies suitable for unstructured data.
@@ 2.3 Human-Centered XAI Evaluations
In practical medical applications, socio-technical gaps may arise between the CDSS explainability factors provided by XAI techniques and end-users’ perceptions of their utility Ackerman (2000). Human-centered evaluations offer methodologies to bridge this gap, aiming to align explanations with user expectations. Such evaluations should hopefully reflect human desired properties of explanations, serving some practical end goal Liao and Xiao (2023).
Human-centered evaluations of XAI systems can be categorized into three levels: (1) application-grounded, (2) human-grounded, and (3) proxy evaluation  Doshi-Velez and Kim (2017). Application-grounded evaluations involve testing real tasks with real experts and should be employed whenever possible because of their fidelity to real-world scenarios. For example, in the medical domain, doctors may be evaluated on specific application tasks. Here, domain experts are required for an effective evaluation. For this reason, application-grounded evaluation may be costly to implement. On the other hand, human-grounded evaluations, use proxy tasks with human evaluations, conducting simplified human-subject experiments. This approach is particularly suitable for assessing general aspects of a system. As such evaluations can be performed with layman participants instead of domain experts, they are not advisable in the medical domain. Finally, proxy evaluations assess proxy tasks without human participation, relying on simulated tasks to provide feedback about a system.
@@
Human-centered evaluation methodologies leverage various techniques, either qualitative or quantitative. In a nutshell, the most widely known are: (1) think-aloud studies, (2) interviews, and (3) surveys. First, in think-aloud studies participants verbalize their thoughts while performing tasks (concurrent) or afterward (retrospective), which can be very useful in qualitatively explaining cognitive processes and identifying usability issues. Second, interviews are a qualitative way to gather detailed information from users about their experiences with a system. They are versatile tools in all stages of design, ranging from open-ended unstructured interviews to tightly controlled structured interviews, with semi-structured interviews offering a valuable compromise. A variant of those is focus groups, in which a small number of participants are gathered to evaluate their opinions about a product or service. Finally, survey methodologies mainly rely on quantitative questionnaires that help researchers gather data on users’ satisfaction, usability issues, and interaction patterns.
@@ 3 Dissecting extant Human-Centered XAI in CDSS
In this section, we present research works that leveraged human-centered XAI to build production-ready CDSS in the medical field. Leveraging the XAI taxonomy outlined before, we discuss papers, highlighting how human-centered evaluations - all relying on application-grounded evaluations - were useful in testing clinical perceptions. In table 1, we provide a summary of the papers discussed. We start discussing CDSS adopting intrinsic interpretable models in Section 3.1. Next, we examine CDSS using post hoc and model agnostic methodologies on top of intrinsic interpretable models in Section 3.2. Finally, we discuss CDSS adopting XAI methodologies in deep learning in Section 3.3, starting with model-agnostic methods, and concluding with model-specific ones.
Paper Model XAI Type Dependency Output Main HCE
Barda et al. (2020) Random Forest SHAP Post Hoc Agnostic Numerical Focus Group
Brennan et al. (2019) Random Forest - Intrinsic Specific Numerical Think Aloud
Ellenrieder et al. (2023) DenseNet IGrads Post hoc Agnostic Visual Think Aloud
Hwang et al. (2022) TRIER (CNN) SG Post Hoc Specific Visual Mixed
Kim et al. (2022) XGBoost SHAP Post Hoc Agnostic Numerical Interviews
Kumarakulasinghe et al. (2020) Random Forest LIME Post Hoc Agnostic Numerical Survey
Kovalchuk et al. (2022) Many SHAP Post Hoc Agnostic Numerical Mixed
Kyrimi et al. (2020) Bayesian Network - Intrinsic Specific Numerical Survey
Matthiesen et al. (2021) Random Forest LIME Post Hoc Agnostic Visual Interviews
Neves et al. (2021) Many Many Post Hoc Agnostic Visual Survey
Pumplun et al. (2023) DenseNet IGrads Post Hoc Agnostic Visual Survey
Rajashekar et al. (2024) Random Forest ICE, ALE Post Hoc Agnostic Numerical Mixed
Sabol et al. (2020) CNN X-CFCMC Post Hoc Agnostic Visual Survey
Singh et al. (2021) Inception-v-3 DeepTaylor Post Hoc Specific Visual Survey
Sivaraman et al. (2023) XGBoost SHAP Post Hoc Agnostic Numerical Think Aloud
Zhang et al. (2024) LSTM Attention Intrinsic Specific Numerical Interviews
Table 1: Summary of papers in alphabetical order adopting XAI methods and Human-Centered Evaluations in medical CDSS. Categorization is executed by defining ML models, XAI methods and their taxonomy (type, dependency, output), and HCE methodologies.
@@ 3.1 CDSS adopting Intrinsic Interpretable Models
Intrinsic interpretable models are easily testable in the medical field for CDSS due to their transparency, as no post-hoc evaluations are necessarily required to perform model evaluations. Examples in this scope are the works from Brennan et al. (2019) and Kyrimi et al. (2020). Brennan et al. (2019) validated the usability of MyRiskSurgery, an analytical CDSS based on generalized linear models and random forests to estimate postoperative risk complications. Through think-aloud studies, a set of physicians evaluated risk complications for a set of 8-10 individual cases in two case studies such as diagnosing acute kidney injury and intensive care unit admission, with the system displaying the top 3 contributing features for each prediction. As a result, physicians experienced notable improvements in their risk assessment, which led to a substantial net improvement in reclassification rates from a no XAI evaluation (+12% and +16%, respectively), also finding the system user-friendly and beneficial.

### v2

@@ Abstract
Explainable AI (XAI) has become a crucial component of Clinical Decision Support Systems (CDSS) to enhance transparency, trust, and clinical adoption. However, while many XAI methods have been proposed, their effectiveness in real-world medical settings remains insufficiently evaluated in practice. This paper presents the first systematic survey of human-centered evaluations (HCE) of XAI methods in CDSS. We categorize existing studies by XAI methodology, evaluation framework, and clinical adoption challenge, synthesizing evidence on how explanations are assessed with clinicians and other stakeholders. Our findings reveal persistent barriers that hinder adoption, such as high cognitive load, misalignment with clinical reasoning, and limited stakeholder engagement. Beyond reviewing prior work, we derive prescriptive insights and propose a stakeholder-centric evaluation framework to guide the design of trustworthy, integrative, and clinically viable CDSS.
@@ 1 Introduction
[…]
Figure 1: The survey exclusively focuses on works at the intersection of Healthcare, Explainable AI, and Human-Centered Evaluations (yellow region).
Artificial Intelligence (AI) is increasingly being integrated into Clinical Decision Support Systems (CDSS) to improve timely diagnosis, optimize treatment plans, and reduce healthcare costs Sendak et al. (2020). Estimates suggest that AI adoption in healthcare could reduce U.S. healthcare expenditures by 5-10%, translating to \$200–\$360 billion in annual savings while maintaining quality and access Sahni et al. (2023). The rapid growth of AI-driven CDSS has also attracted significant investment, with venture capital funding in AI in health exceeding \$11 billion in 2024 Goldsack and Overgaard (2024). These trends indicate that the role of AI in medical decision-making will continue to expand Lorenzini and others (2023).
However, a key challenge accompanying this expansion is the necessity for trust and transparency in AI-driven medical decisions Qin and others (2024). Despite advancements in deep learning and other black-box AI models, clinicians often hesitate to adopt AI-generated recommendations Longoni et al. (2019)Gaczek and others (2023), often due to their lack of interpretability Hou and others (2024). Trust in CDSS is critical because errors in automated decision-making can have life-threatening consequences Zhang and Zhang (2023). In response, Explainable AI (XAI) has emerged as a solution to improve model transparency, thereby enabling medical professionals to better understand why an AI system makes certain predictions Loh et al. (2022). Additionally, regulatory frameworks, such as the General Data Protection Regulation (GDPR), emphasize the need for explainability in automated decision-making, thereby necessitating interpretable AI solutions in healthcare.
@@
To address this gap, this paper provides the first systematic survey of human-centered evaluations (HCE) of XAI methods in CDSS. Unlike prior surveys on XAI in healthcare (e.g., Antoniadi et al. (2021); Loh et al. (2022); Kim and others (2024)), which emphasize algorithmic innovations or cross-domain interpretability frameworks, our focus is on how explanations are actually assessed with clinicians and other stakeholders in practice. Specifically, we: (1) categorize existing works based on XAI methodologies, evaluation frameworks, and clinical adoption challenges, offering a structured synthesis of the field; (2) identify key gaps in evaluation practices, highlighting inconsistencies and challenges such as high cognitive load, misalignment with clinical knowledge, and limited stakeholder engagement; (3) introduce a conceptual framework to bridge the socio-technical gap in AI-driven CDSS by aligning XAI evaluation methods with diverse stakeholder needs (e.g., clinicians, administrators, AI developers); and (4) propose a stakeholder-centric approach for CDSS development, advocating that CDSS be designed as human-augmentation tools rather than replacements.
To theoretically ground the misalignment between human expectations and technological capabilities, we draw on Ackerman’s socio-technical gap theory, which highlights the challenges such discrepancies create Ackerman (2000). This analysis informs our proposed framework, emphasizing stakeholder-driven evaluations to ensure CDSS solutions are clinically relevant, interpretable, and effectively integrated into healthcare workflows. By aligning technical advancements in XAI with rigorous human-centered evaluations, we aim to steer research toward developing trustworthy, deployable CDSS solutions.
@@ 2 Background
In this section, we provide a brief overview on CDSS; showcase a taxonomy on XAI, detailing criteria based on type, dependency, scope and output; review related work on XAI in healthcare, pointing out a lack of human-centered evaluation; and, in turn, review how such evaluations can be categorized and leveraged. Figure 1 shows the focus of this review.
@@ 2.1 Clinical Decision Support Systems (CDSS)
CDSS have long been integral to medical decision-making, assisting clinicians in enhancing patient outcomes. Initially emerging in the late 1950s with rule-based systems Ledley and Lusted (1959), CDSS evolved significantly with advancements in AI. Modern CDSS now leverage sophisticated machine learning algorithms to process vast datasets efficiently. Concurrently, the rise of XAI has enabled the integration of interpretable solutions, essential for fostering trust and promoting the adoption of CDSS in clinical practice Ghassemi et al. (2021).
Operationally, CDSS utilize patient data to generate specific treatment recommendations Musen et al. (2021). The back-end employs machine learning algorithms trained for specific tasks, processing data from Electronic Health Records (EHRs). EHRs include both structured data (e.g., tabular data) and/or unstructured data (e.g., text, images), representing comprehensive digital versions of patients’ medical histories, including diagnoses, treatments, and test results. System performance is evaluated using machine learning metrics pertinent to the specific task. Synchronously, the front-end runs a user interface (UI) connected to the back-end that clinical practitioners are supposed to interact with. Such UI displays patient data, alerts, recommendations (sometimes with explanations), and decision-support reports in an accessible manner to facilitate clinical decision-making. To this end, human-centered evaluations are conducted to assess the overall effectiveness and efficiency of the system in the specific task. However, researchers have identified that the absence of such evaluations is a primary factor in the lack of adoption of AI-based CDSS solutions Musen et al. (2021), particularly due to insufficient consideration of clinicians’ workflows and the collaborative nature of clinical practice Kohli and Jha (2018); Musen et al. (2021); Wears and Berg (2005); Horsky et al. (2012).
2.2 A Taxonomy of XAI Methods
Many taxonomies have been proposed to categorize XAI methodologies (e.g., see Speith (2022)). In this paper, we adopt the conceptual approach based on the following macro criteria (for a more comprehensive review, please refer to Molnar (2022)):
@@ 2.2.4 Output: Numerical or Visual
Visual output methods are particularly suitable for deep learning models operating with unstructured data such as images. Visualization techniques, such as saliency maps, class activation maps (CAM), and GradCAM, enable the interpretation of feature importance in CNNs, highlighting regions of interest in input images. Parallely, this also extends to token importance visualisation for language models. Beyond deep learning, structured data can also benefit from visual representations, including partial dependence plots, feature importance bar charts, and decision boundary plots. However, we treat those methodologies as numerical. Methods: all methodologies suitable for unstructured data.
2.3 Related Work on XAI Methods in Healthcare
We briefly overview extant literature using the XAI methods mentioned in the taxonomy previously discussed.
SHAP emerged as one of the most widely adopted XAI technique, utilized for identifying critical clinical features in disease prediction and patient outcome assessment Liu and others (2022); Shi and others (2022); Chen et al. (2021). The technique proved especially valuable for hospital management applications, including mortality prediction in intensive care units, hospital readmission risk assessment, and surgical complication forecasting, with clinical feature importance analysis revealing consistent patterns where age, laboratory values, and vital signs were frequently identified as primary predictors Nguyen and others (2021); Zeng and others (2021); Zhang and others (2021).
Next, GradCAM ranked first for visual explanation tasks, focusing on medical imaging interpretation Figueroa and others (2022); Xu and others (2021); Chetoui et al. (2021). Also, this technique proved particularly effective for COVID-19 detection from chest X-rays, with studies demonstrating its utility in highlighting lung discriminative regions for pandemic-related diagnosis Shi et al. (2021); Ozturk et al. (2020). GradCAM implementations primarily utilized deep learning architectures such as ResNet and VGG, extending successfully beyond radiological applications to ophthalmological imaging, dermatological lesion analysis, and histopathological examination Yoo et al. (2021); Thakoor et al. (2021).
Other XAI techniques have been applied in more specialized contexts, with LIME adopted for single-instance explanations, Layer-wise Relevance Propagation (LRP) utilized for heatmap generation, and various other methods including fuzzy classifiers employed for specific clinical tasks Binder and others (2021); Sabol et al. (2020). Finally, XAI applications in clinical NLP showed significant potential for medical coding automation and mental health assessment, with attention mechanisms achieving AUROC scores of 0.92 and 0.90 for medical code prediction tasks Dong et al. (2021); Hu et al. (2021).
However, despite a general positive trend of XAI adoption for clinical usage, those systems fail to consider human feedback as an evaluation principle, leaving socio-technical gaps between clinical practitioners and technological solutions, limiting CDSS effectiveness for an effective patient-centric care. Motivated by that, in the next subsection, we provide a review on human-centered XAI evaluations.
@@ 2.4 Human-Centered XAI Evaluations
In practical medical applications, socio-technical gaps may arise between the CDSS explainability factors provided by XAI techniques and end-users’ perceptions of their utility Ackerman (2000). Human-centered evaluations offer methodologies to bridge this gap, aiming to align explanations with user expectations. Such evaluations should hopefully reflect human desired properties of explanations, serving some practical end goal Liao and Xiao (2023). Human-centered evaluations of XAI systems can be categorized into three levels: (1) application-grounded, (2) human-grounded, and (3) proxy evaluation Doshi-Velez and Kim (2017).
Application-grounded evaluations involve testing real tasks with real experts and should be employed whenever possible because of their fidelity to real-world scenarios. For example, in the medical domain, doctors may be evaluated on specific application tasks. Here, domain experts are required for an effective evaluation. For this reason, application-grounded evaluation may be costly to implement. On the other hand, human-grounded evaluations use proxy tasks with human evaluations, conducting simplified human-subject experiments. This approach is particularly suitable for assessing general aspects of a system. As such evaluations can be performed with layman participants instead of domain experts, they are not advisable in the medical domain. Finally, proxy evaluations assess proxy tasks without human participation, relying on simulated tasks to provide feedback about a system.
@@
Human-centered evaluation methodologies leverage various techniques, either qualitative or quantitative. In a nutshell, the most widely known are: (1) think-aloud studies, (2) interviews, and (3) surveys. First, in think-aloud studies participants verbalize their thoughts while performing tasks (concurrent) or afterward (retrospective), which can be very useful in qualitatively explaining cognitive processes and identifying usability issues. Second, interviews are a qualitative way to gather detailed information from users about their experiences with a system. They are versatile tools in all stages of design, ranging from open-ended unstructured interviews to tightly controlled structured interviews, with semi-structured interviews offering a valuable compromise. A variant of those is focus groups, in which a small number of participants are gathered to evaluate their opinions about a product or service. Finally, survey methodologies mainly rely on quantitative questionnaires to gather data on users’ satisfaction, usability issues, and interaction patterns.
@@ 3 Dissecting extant Human-Centered XAI in CDSS
We survey AI on healthcare HCI conference sessions and healthcare journals to form a database of possible candidate research papers. In parallel, we query research search engines such as Google Scholar, ACM Digital Library and PubMed with a set of keywords, including Clinical Decision Support System (CDSS) and Explainable AI (XAI) to further enrich the database. Our inclusion criteria comprise: (1) top-tier conference venue or Q1 journal; (2) publication date after 2017, motivated by the DARPA XAI Program announced in May 2017 Gunning and Aha (2019), which fostered general innovation in XAI, and the subsequent development and wide adoption of AI black-box models and XAI tools such as LIME and SHAP to explain their predictions; (3) development of a concrete CDSS, usable and integrable into existing clinical practice (this automatically excludes simulation tools or akin); (4) strong reliance on at least one HCI method adopted with clinical practitioners, gathering their feedback on the CDSS. We manually review each paper retrieved to verify the latter. As a result, we are left with 17 papers encompassing the three macro-categories as shown in Figure 1 and the inclusion criteria. In table 1, we provide a summary of the papers retrieved.
We firstly start discussing CDSS adopting intrinsic interpretable models in Section 3.1. Next, we examine CDSS using post hoc and model agnostic methodologies on top of intrinsic interpretable models in Section 3.2. Finally, we discuss CDSS adopting XAI methodologies in deep learning in Section 3.3, starting with model-agnostic methods, and concluding with model-specific ones.
Paper ML Model XAI Type Dependency Output HCE #Subjects
Barda et al. (2020) Random Forest SHAP Post Hoc Agnostic Numerical Focus Group 8
Brennan et al. (2019) Random Forest - Intrinsic Specific Numerical Think Aloud 20
Ellenrieder et al. (2023) DenseNet IGrads Post hoc Agnostic Visual Think Aloud 42
Hwang et al. (2022) TRIER (CNN) SG Post Hoc Specific Visual Mixed 9
Kim et al. (2022) XGBoost SHAP Post Hoc Agnostic Numerical Interviews 36
Kumar. et al. (2020) Random Forest LIME Post Hoc Agnostic Numerical Survey 10
Kovalchuk et al. (2022) Many SHAP Post Hoc Agnostic Numerical Mixed 19
Kyrimi et al. (2020) Bayesian Network - Intrinsic Specific Numerical Survey 16
Matthiesen et al. (2021) Random Forest LIME Post Hoc Agnostic Visual Interviews 7
Neves et al. (2021) Many Many Post Hoc Agnostic Visual Survey 3
Pumplun et al. (2023) DenseNet IGrads Post Hoc Agnostic Visual Survey 45
Rajashekar et al. (2024) Random Forest ICE, ALE Post Hoc Agnostic Numerical Mixed 31
Sabol et al. (2020) CNN X-CFCMC Post Hoc Agnostic Visual Survey 14
Singh et al. (2021) Inception-v-3 DeepTaylor Post Hoc Specific Visual Survey 14
Sivaraman et al. (2023) XGBoost SHAP Post Hoc Agnostic Numerical Think Aloud 24
Zhang et al. (2024) LSTM Attention Intrinsic Specific Numerical Interviews 6
Bienefeld et al. (2023) Many SHAP Post Hoc Agnostic Numerical Many 95
Table 1: Summary of papers adopting XAI methods and Human-Centered Evaluations in medical CDSS. Categorization is based on ML models, XAI methods, their taxonomy (type, dependency, output), and HCE methodologies.
@@ 3.1 CDSS adopting Intrinsic Interpretable Models
Intrinsic interpretable models are easily testable in the medical field for CDSS due to their transparency, as no post-hoc evaluations are necessarily required to perform model evaluations. Examples in this scope are the works from Brennan et al. (2019) and Kyrimi et al. (2020). Brennan et al. (2019) validated the usability of MyRiskSurgery, an analytical CDSS based on generalized linear models and random forests to estimate postoperative risk complications. Through think-aloud studies, a set of physicians evaluated risk complications for a set of 8-10 individual cases in two case studies such as diagnosing acute kidney injury and intensive care unit admission, with the system displaying the top 3 contributing features for each prediction. As a result, physicians experienced notable improvements in their risk assessment, which led to a substantial net improvement in reclassification rates from a no XAI evaluation (+12% and +16%, respectively), also finding the system and beneficial.